Given a sample size of 20, what are the critical values X 2/L and X 2/R for a 95% confidence level?
Answer #1

df = n-1 = 20 - 1 = 19

Chi square critical values at 0.05 level with 19 df are

\chi^2L = 8.907 , \chi^2R = 32.852